On a cold winter day, Pelle spills out 1.0 dl of water with a temperature of 20 ° C. Two minutes later, all the water has frozen. 

Determine the average power with which the water transfers heat to the surroundings during the two minutes.
1
Expert's answer
2021-03-14T19:17:06-0400
"Pt=mL+mc\Delta T=m(L+c\Delta T)\\
(2\cdot 60 )P=0.1(334000+4200\cdot20)\\P=348\ W"